Presentation on theme: "Module Production T. Bowcock. Substrates Twist problem –Further understanding of the warp/weft and layup problem Chirality of the twist –Looks like it."— Presentation transcript:
Module Production T. Bowcock
Substrates Twist problem –Further understanding of the warp/weft and layup problem Chirality of the twist –Looks like it can be controlled It looks like it is a function of the 45°layer. –The current layup is good ~0.5° (Limit about 3°) Pressure identified as part of problem –Reproduce twist at Liverpool by increasing pressure >90psi –Note nominal maximum pressure ~300psi Determine the CF/Epoxy ratios –For our modules ~ 9.9g/Weight (module) –Should be approx 60-68% CF by weight. PAC still quotes at 10/week.
K-05 (Re)-Processing almost complete –Last batch were untestable through overetching Circuits should be ready to laminate next week at Stevenage
Pitch Adaptors All pitch adaptors sent back to CERN for plasma cleaning Current ones not testable –2 controls done in old pas –Damage/deposit on needle
Sensors 8 arrived at Liverpool –Currents ~uA and 0<0.5% percent strips bad. –One with no bad strips Perfect cutting 2 returned to Micron for remeasurement (3 others high in dB) –One recleaned at Micron did not change current
Sensor Testing Problem On new batch of sensors on some detectors measure substantially higher current than micron Complete halt and check for damage to sensors –Inspection of old sensors –Running tests with mechanicals –Tests with quartz Repeat of tests using chuck/kapton –Currents confirmed
Sensor Testing Limited evidence of microscopic liquid from vacuum holes Hypothesis that the jigs were contaminating the sensors (Noryl) 2 nd and 3 rd tier of cleaning of jigs Tests underway 8 Pairs of new (cleaner) jigs under construction for Liverpool and Micron –Acetal rather than Noryl
Bonding & Jigs Now have 4 double sided bonding jigs –1 for backend –2 for frontend –1 for laser system 2 handling frames created (expect about 10 at the end)
Module Production Feet and Bases still in production Paddle Moulds (3 off) still to be made
Testing Equipment Laser rig plate finalized and submitted Thermograph equipment designed (to be reviewed) Gluing equipment expect to start the metrology and programming next week
Summary K-05 circuits close to delivery Twist story still approx 2 weeks from resolution –Living with the twist? Pitch adaptor problems again. Need cleaning? Sensor testing issues uncovered –Micron results look OK All mechanics except the thermograph system and the final paddle-hybrid assembly (twist) ready by July.
Timeline No prototype modules will be completed until end Aug 05 based on current status of K-05 and twist (guesstimate) –4 double sided modules Look for real production start ~ Oct 05